OpenStudy (russianboss1):

The length of a rectangle is 8 mm longer than its width. Its perimeter is more than 36 mm. Let w equal the width of the rectangle. Write an inequality to model this information. A) w+ (w+8)>36 B)2(w+8)+2w>36 C)2(w+8)+2w<36 D)(w+8)+w<36 Fan and Medal

OpenStudy (anonymous):

If the perimeter formula is this: \[P = 2L + 2W\] And, our givens are: \[L = W + 8\] \[2L+2W>36\] What's our inequality? Can you connect the dots? @RussianBoss1 Hint: Substitute L

OpenStudy (anonymous):



OpenStudy (russianboss1):

@AloneS @robtobey @IrishBoy123 @BruhObamaCare @FortyTheRapper @Crazy_questions @Ghostgate

OpenStudy (anonymous):

Really? @RussianBoss1 BIG HINT: \[~~~L = (W + 8)\]\[~~\downarrow\]\[~2L+2W>36\]\[~~\downarrow\]\[2(?) + 2W > 36\] Please take a look at your choices.